600 500 400 300 200 100 0-10 11-21 22-32 33-43 44-54 55-65 66-76 77-87 88+ Salary (S thousands) Number of employees 8

a.) What is the approximate number of employees who earn between $33,000 and $43,000 annually?

b.) There are approximately 2,700 employees of ACME Corporation. About what percentage of employees earn over $76,999?

c.) The largest percentage of employees earn between $44,000-$54,000. Which 2 salary ranges have roughly an equal amount of employees?
Answer #1

Answer

(A) using the given histogram, check the number of employees on the y axis who earn between $33,000 and $43,000 annually

we get, n = 400

(B) total number of employees n = 2700

number of employees above $76,999 = 350+100 = 450

Required % = (number of employees above $76,999 /total)*100

= (450/2700)*100

= 16.67%

(C) using the given histogram, check the number of employees on the y axis who has exactly same height of histogram column

so, we get

22 to 32 and 66 to 76 range have roughly an equal amount of employees